On January 1, 2018, Badger Inc. adopted the dollar-value LIFO method. The inventory cost on this date was $100,700. The ending inventory, valued at year-end costs, and the relative cost index for each of the next three years is below:

Year-end Ending inventory at
year-end costs
Cost Index
2018 $ 128,205 1.05
2019 146,080 1.10
2020 156,540 1.20

In determining the inventory balance should Badger report in its 12/31/2019 balance sheet:
Answer #1

Calculate following

Base year price = 128205/1.05 = 122100

Bae year price 2019 = 146080/1.1 = 132800

Inventory balance 2019 as per Dollar value LIFO method = 100700+21400*1.05+10700*1.1 = 134940

In determining the inventory balance should Badger report in its 12/31/2019 balance sheet: $134940